Calculate the time that is required to electroplate a 1.0 mm thick silver layer on a 2.0 × 2.0 cm2 surface when a current of 1.0 A is allowed to pass through the electrolyte.
Answer #1

Density of silver is 10.5 g/cm3.

First calculate the area which is electroplate by silver

Thickness = 1.0 mm= 0.1 cm

Volume = thickness * area

= 2.0 × 2.0 cm2 × 0.1 cm

= 0.4 cm^3

Density = mass / volume

Mass = density * volume

= 10.5 g/cm3 *0.4 cm^3

= 4.2 g


Applying to silver E = Z × 96500
Z = 108/96500 g
Let the current be passed for t seconds.
We know that
W = Z × I × t
So, 4.2 = 108/96500 x 1.0 x t
or t = (4.2 × 96500)/(108×1.0)

=3752.77 second

= 62.55 min

= 1.04 hr